# Adversarial Score Distillation: When score distillation meets GAN

### v1

@@ Abstract
Existing score distillation methods are sensitive to classifier-free guidance (CFG) scale: manifested as over-smoothness or instability at small CFG scales, while over-saturation at large ones. To explain and analyze these issues, we revisit the derivation of Score Distillation Sampling (SDS) and decipher existing score distillation with the Wasserstein Generative Adversarial Network (WGAN) paradigm. With the WGAN paradigm, we find that existing score distillation either employs a fixed sub-optimal discriminator or conducts incomplete discriminator optimization, resulting in the scale-sensitive issue. We propose the Adversarial Score Distillation (ASD), which maintains an optimizable discriminator and updates it using the complete optimization objective. Experiments show that the proposed ASD performs favorably in 2D distillation and text-to-3D tasks against existing methods. Furthermore, to explore the generalization ability of our WGAN paradigm, we extend ASD to the image editing task, which achieves competitive results. The project page and code are at this link.
@@ 1 Introduction
However, as a milestone in score distillation methods, Score Distillation Sampling (SDS) [43] is very sensitive to the classifier-free guidance (CFG) [19] scale. Figure 1 takes 2D score distillation as an example. When setting a small CFG scale, SDS tends to obtain over-smoothing results, whereas when setting a large one, the generated images become over-saturated. Recently, Variational Score Distillation (VSD) [55] alleviates the over-smoothing problem at small CFG scales, but we observe that VSD is a bit unstable during the distillation process at small CFG scales, as shown in Figure 1. These phenomena make us conjecture that there are still some unrevealed methodological issues behind SDS and VSD.
@@
Figure 2: Examples generated by ASD in 2D distillation, image editing, and text-to-3D tasks. Stable Diffusion [51, 50] is used as the pretrained diffusion model.
@@ 3 Preliminaries
Generative Adversarial Networks. Generative Adversarial Networks (GANs), first introduced in  [13], can be formulated by a zero-sum competition between a discriminator network $D$ and a generator network $G$. From an intuitive understanding, the discriminator is trained to classify images as fake or real, while the generator is trained to generate images that fool the discriminator. The original form of GAN [13] essentially minimizes the Jensen-Shannon (JS) divergence between the distributions of real samples and generated samples, which gets a gradient close to zero when two distributions are very different. The Wasserstein GAN (WGAN) [3, 56] reformulates the original GAN with the Wasserstein distance to handle the zero-gradient issue, which can be written as
@@ 4 Adversarial Score Distillation
where $p(y|x_{t})$ denotes the probability of the given noisy sample $x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon$ to be categorized as the real samples described by prompt $y$, $p(\phi|x_{t})$ represents the probability of $x_{t}$ to be categorized as the fake samples described by prompt $\phi$. When $x_{t}$ is close to samples described by prompt $y$, this $D(x_{t};y)$ will give high confidence. Otherwise, it will get low confidence when $x_{t}$ is close to samples described by prompt $\phi$. Based on WGAN [3, 56], we have the generator loss as
@@
where $\tilde{x}$ comes from a mixture distribution $\tilde{\mu}$, so it can be replaced by a combination of $x_{t}^{r}$ and $x_{t}^{g}$. $\eta$ and $\gamma$ are adjustable hyperparameters, to keep the penalty term positive, they are subject to $\gamma\geq-\eta\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon_{x_{t}^{r};\phi,t}\rVert_{2}^{2}/\lVert\epsilon_{x_{t}^{g};y,t}-\epsilon_{x_{t}^{g};\phi,t}\rVert_{2}^{2}$. In practice, we find $\eta=1$, $\gamma\in[-1,0)$ works for most cases. See supplementary materials for the derivation details of Eq. (9). In SDS [43], this discriminator optimization is completely ignored, as there is no optimizable $\phi$. VSD [55] uses the LoRA or simple UNet branch to implement $\phi$, and updates it using $\nabla_{\phi}\lVert\epsilon_{x_{t}^{g};\phi,t}-\epsilon\rVert_{2}^{2}$, which is just a part of Eq. (9).
Figure 3: Workflow of ASD. Green lines show the pipeline of generator optimization. Orange lines show the pipeline of discriminator optimization. The avatar of NeRF is adapted from [39].
@@
$\mathcal{L}^{\prime}_{D}$ for text-conditioned distillation. When $\eta=1$, based on triangle inequality, we have $\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon_{x_{t}^{r};\phi,t}\rVert_{2}^{2}\leq\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon\rVert_{2}^{2}+\lVert\epsilon_{x_{t}^{r};\phi,t}-\epsilon\rVert_{2}^{2}$. Thus, we can rewrite Eq. (9) as
@@
SDS (CFG = 7.5) ${}^{\dagger}$ (CFG = 7.5)
Figure 4: 2D score distillation results with the prompts “hamburger” and “a monster truck”. ${}^{\dagger}$ denotes that the LoRA branch is updated with 50 steps per iteration, resulting in over-smoothing images similar to SDS.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 5 Comparison and Discussion
(a) 2D score distillation progress.
Optimization Steps
@@
(b) Text-to-3D NeRF distillation progress.
Figure 7: Stability comparisons of the distillation progress between VSD and SDS. The prompts are “a realistic happy dog playing in the grass” and “a pineapple” for 2D distillation and the text-to-3D, respectively. As can be seen, ASD constantly keeps the main object and image / NeRF structures unchanged.
Disscussion of $\gamma$. When $\gamma=0$, Eq (10) becomes VSD, which can be considered as the WGAN [3, 56] discriminator loss with fixed penalty term for Lipschitz constraint. For $\gamma<0$, we decrease the weight of the penalty term, which can enhance the ability of the discriminator. For $\gamma>0$, we increase the weight of the penalty term. Specifically, when $\gamma\gg 0$, the penalty term will dominant Eq (10), which leads to $\epsilon_{x_{t}^{g},y,t}\approx\epsilon_{x_{t}^{g},\phi,t}$. Then the discriminator fails to recognize fake samples, and the gradient for generators (6) mainly comes from the implicit classifier term. Only using the classifier term can be considered as a special case of our paradigm in this situation.
@@ 6 Experiments
We first evaluate ASD in 2D score distillation and text-to-3D tasks in terms of quality, stability, and diversity. Then we extend ASD to image editing tasks and obtain some competitive results.
@@
Optimization Steps
DDS Refer to caption Refer to caption Refer to caption
ASD Refer to caption Refer to caption Refer to caption
a cartoon wolf $\rightarrow$ a cartoon wolf
DDS Refer to caption Refer to caption Refer to caption
ASD Refer to caption Refer to caption Refer to caption
a tiger $\rightarrow$ a cartoon tiger
DDS Refer to caption Refer to caption Refer to caption
ASD Refer to caption Refer to caption Refer to caption
a squirrel is eating a piece $\rightarrow$ a panda is eating a piece
of food on the ground           of food on the ground
Figure 9: Comparisons between DDS and ASD on 2D image editing tasks. We test our ASD in three different situations, simplification, refinement, and replacement of image editing tasks.
@@ 7 Limitation
Despite showing favorable capabilities in terms of distillation quality, stability, and diversity, the speed of ASD is similar to VSD. We find that sharing the same added noise between generator optimization and discriminator optimization can speed up distillation by 30 $\sim$ 50%, with a little quality sacrifice. However, distilling a 3D NeRF still takes more than 5 hours on a single A800 GPU. Using faster PEFT [20, 34, 32, 11, 30] methods to update the discriminator or designing forward-only methods to approximate $\epsilon_{x_{t};\phi,t}$ may be promising solutions to this limitation.
@@ 8 Conclusion
In this paper, we proposed the Adversarial Score Distillation (ASD) under the WGAN paradigm. Compared to existing score distillation methods, our ASD exploits a complete discriminator loss with a principled discussion, which alleviates the CFG-sensitive issue and leads to impressive performances.
@@ Appendix A Quantitative Evaluations and User Studies for 3D Generation
[Uncaptioned image] [Uncaptioned image]
@@ Appendix B Derivation of Adversarial Score Distillation Loss and Gradients
where we use $-\log p(x_{0})\approx\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t},\phi,t}-\epsilon\rVert_{2}^{2}]$, according to [18, 28]. And $\tilde{x}$ comes from a mixture distribution $\tilde{\mu}$, so it can be replaced by a combination of $x_{t}^{r}$ and $x_{t}^{g}$. $\eta$ and $\gamma$ are adjustable hyperparameters, to keep the penalty term positive, they are subject to $\gamma\geq-\eta\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon_{x_{t}^{r};\phi,t}\rVert_{2}^{2}/\lVert\epsilon_{x_{t}^{g};y,t}-\epsilon_{x_{t}^{g};\phi,t}\rVert_{2}^{2}$. In practice, we find $\eta=1$, $\gamma\in[-1,0)$ works for most cases.

### v2

@@ Abstract
Existing score distillation methods are sensitive to classifier-free guidance (CFG) scale, manifested as over-smoothness or instability at small CFG scales, while over-saturation at large ones. To explain and analyze these issues, we revisit the derivation of Score Distillation Sampling (SDS) and decipher existing score distillation with the Wasserstein Generative Adversarial Network (WGAN) paradigm. With the WGAN paradigm, we find that existing score distillation either employs a fixed sub-optimal discriminator or conducts incomplete discriminator optimization, resulting in the scale-sensitive issue. We propose the Adversarial Score Distillation (ASD), which maintains an optimizable discriminator and updates it using the complete optimization objective. Experiments show that the proposed ASD performs favorably in 2D distillation and text-to-3D tasks against existing methods. Furthermore, to explore the generalization ability of our paradigm, we extend ASD to the image editing task, which achieves competitive results. The project page and code are at this link.
@@ 1 Introduction
However, as a milestone in score distillation methods, Score Distillation Sampling (SDS) [44] is very sensitive to the classifier-free guidance (CFG) [18] scale. Figure 1 takes 2D score distillation as an example. When setting a small CFG scale, SDS tends to obtain over-smoothing results, whereas when setting a large one, the generated images become over-saturated. Recently, Variational Score Distillation (VSD) [56] alleviates the over-smoothing problem at small CFG scales, but we observe that VSD is a bit unstable during the distillation progress at small CFG scales, as shown in Figure 1. These phenomena make us conjecture that there are still some unrevealed methodological issues behind SDS and VSD.
@@
Figure 2: Examples generated by ASD in 2D distillation, image editing, and text-to-3D tasks. Stable Diffusion [52, 51] is used as the pretrained diffusion model. For more results please refer to our project page.
@@ 3 Preliminaries
Generative Adversarial Networks. Generative Adversarial Networks (GANs), first introduced in  [13], can be formulated by a zero-sum competition between a discriminator network $D$ and a generator network $G$. The original form of GAN [13] essentially minimizes the Jensen-Shannon (JS) divergence between the distributions of real samples and generated samples, which gets a gradient close to zero when two distributions are very different. The Wasserstein GAN (WGAN) [3, 58] reformulates the original GAN with the Wasserstein distance to handle the zero-gradient issue, which can be written as
@@ 4 Adversarial Score Distillation
where $p(y|x_{t})$ denotes the probability of the given noisy sample $x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon$ to be categorized as the samples described by prompt $y$, $p(\phi|x_{t})$ represents the probability of $x_{t}$ to be categorized as the samples described by prompt $\phi$. Here, we consider $\phi$ as the prompt for fake samples. When $x_{t}$ is close to samples described by prompt $y$, this $D(x_{t};y)$ will give high confidence. Otherwise, it will get low confidence when $x_{t}$ is close to samples described by prompt $\phi$. Based on WGAN [3, 58], we have the generator loss as
@@
where $\tilde{x}$ comes from a mixture distribution $\tilde{\mu}$, so it can be replaced by a combination of $x_{t}^{r}$ and $x_{t}^{g}$. $\eta$ and $\gamma$ are adjustable hyperparameters, to keep the penalty term positive, they are subject to $\gamma\geq-\eta\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon_{x_{t}^{r};\phi,t}\rVert_{2}^{2}/\lVert\epsilon_{x_{t}^{g};y,t}-\epsilon_{x_{t}^{g};\phi,t}\rVert_{2}^{2}$. In practice, we find $\eta=1/2$, $\gamma\in[-1,0)$ works for most cases. See supplementary materials for the derivation details of Eq. (9). In SDS [44], this discriminator optimization is completely ignored, as there is no optimizable $\phi$. VSD [56] uses the LoRA or simple UNet branch to implement $\phi$, and updates it using $\nabla_{\phi}\lVert\epsilon_{x_{t}^{g};\phi,t}-\epsilon\rVert_{2}^{2}$, which is just a part of Eq. (9).
Figure 3: Workflow of ASD. Green lines show the pipeline of generator optimization. Orange lines show the pipeline of discriminator optimization. The avatar of NeRF is adapted from [40]. See the supplementary for the algorithm description.
@@
$\mathcal{L}^{\prime}_{D}$ for text-conditioned distillation. When $\eta={1}/{2}$, based on triangle and Cauchy–Schwarz inequality, we have $\frac{1}{2}\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon_{x_{t}^{r};\phi,t}\rVert_{2}^{2}\leq\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon\rVert_{2}^{2}+\lVert\epsilon_{x_{t}^{r};\phi,t}-\epsilon\rVert_{2}^{2}$. Thus, we can rewrite Eq. (9) as
@@
SDS (CFG = 7.5) VSD (CFG = 7.5)
Figure 4: 2D score distillation results with the prompts “hamburger” and “a monster truck”. VSD denotes that the LoRA branch is updated with 50 steps per iteration, resulting in over-smoothing images similar to SDS.
@@ 5 Comparison and Discussion
Disscussion of $\gamma$. When $\gamma=0$, Eq (10) becomes VSD, which can be considered as the WGAN [3, 58] discriminator loss with fixed penalty term for Lipschitz constraint. For $\gamma<0$, we decrease the weight of the penalty term, which can enhance the ability of the discriminator. For $\gamma>0$, we
@@
increase the weight of the penalty term. Specifically, when $\gamma\gg 0$, the penalty term will dominant Eq (10), which leads to $\epsilon_{x_{t}^{g},y,t}\approx\epsilon_{x_{t}^{g},\phi,t}$. Then the discriminator fails to recognize fake samples, and the gradient for generators (6) mainly comes from the implicit classifier term. Only using the classifier term can be considered as a special case of our paradigm in this situation.
@@
a realistic happy dog playing in the grass
@@
a pineapple
Figure 7: Stability comparisons of the distillation progress between VSD and SDS in both 2D and 3D. As can be seen, ASD constantly keeps the main object and image / NeRF structures unchanged.
@@ 6 Experiments
DDS Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
ASD Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
a cartoon wolf $\rightarrow$ a cartoon wolf a tiger $\rightarrow$ a cartoon tiger a squirrel is * $\rightarrow$ a panda is *
* eating a piece of food on the ground
Figure 9: Comparisons between DDS and ASD on 2D image editing tasks. We test our ASD in three image editing situations: simplification, refinement, and replacement. For each case, we use three images to show the optimization progress.
@@
Quantitative Results. Table 1 presents quantitative evaluations of CLIP scores and user studies for 3D generation tasks, demonstrating that ASD performs favorably against SDS and VSD. See the supplementary for more discussion about CLIP scores.
Method CLIP Score
CLIP B/32$\uparrow$ CLIP B/16$\uparrow$ CLIP L/14$\uparrow$
DreamFusion (SDS) 0.3282 0.3290 0.2889
ProlificDreamer (VSD) 0.3203 0.3258 0.2789
Ours (ASD) 0.3314 0.3376 0.2854
Method User Study
Details(%) Semantic Alignment (%) Overall Quality (%)
ProlificDreamer (VSD) 35.3 45.7 34.5
Ours (ASD) 64.7 54.3 65.5
Table 1: Quantitative Results. CLIP scores are the cosine similarity between multiple randomly rendered views of the 3D model and the given text prompt. User evaluation is conducted between VSD and ASD from three perspectives.
@@ 7 Limitation
Despite showing favorable capabilities in terms of distillation quality, stability, and diversity, the speed of ASD is similar to VSD. We find that sharing the same added noise between generator optimization and discriminator optimization can speed up distillation by 30 $\sim$ 50%, with a little quality sacrifice. However, distilling a 3D NeRF still takes more than 5 hours on a single A800 GPU. Using faster 3D representation [26], or faster PEFT [20, 35, 33, 11, 31] methods to update the discriminator, or designing forward-only methods to approximate $\epsilon_{x_{t};\phi,t}$ may be promising solutions to this limitation.
@@ 8 Conclusion
In this paper, we proposed the Adversarial Score Distillation (ASD) under the WGAN paradigm. Compared to existing methods, our ASD exploits a complete discriminator loss with a principled discussion, which alleviates the CFG-sensitive issue and leads to impressive performances.
@@
9 Acknowledgements
We would like to express our gratitude to all the individuals who provided assistance in this work: Chaohui Yu, Fan Wang. We also thank the authors of all the open-source codes and models that we used.
@@ Appendix A Quantitative Evaluations and User Studies for 3D Generation
Refer to caption Refer to caption
@@ Appendix B Derivation of Adversarial Score Distillation Loss and Gradients
where we use $-\log p(x_{0})\approx\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t},\phi,t}-\epsilon\rVert_{2}^{2}]$, according to [19, 29]. $\tilde{x}$ comes from a mixture distribution $\tilde{\mu}$, so we can sampling both real or fake data points [58] which means $\tilde{x}_{t}$ can be replaced by a combination of $x_{t}^{r}$ and $x_{t}^{g}$. $\eta$ and $\gamma$ are adjustable hyperparameters, to keep the penalty term positive, they are subject to $\gamma\geq-\eta\lVert\epsilon_{x_{t}^{r};y,t}-\epsilon_{x_{t}^{r};\phi,t}\rVert_{2}^{2}/\lVert\epsilon_{x_{t}^{g};y,t}-\epsilon_{x_{t}^{g};\phi,t}\rVert_{2}^{2}$. However, in Eq. (16), the noisy real sample $x_{t}^{r}$ is unavailable for text-conditioned cases. For these cases, we can use the upper bound of $\mathcal{L}_{D}$, (i.e. $\mathcal{L}^{\prime}_{D}$), which does not contain $x_{t}^{r}$ terms. When $\eta=1/2$, based on triangle and Cauchy–Schwarz inequality, we have:
$\nabla_{\phi}\mathcal{L}^{\prime}_{D}=\nabla_{\phi}\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t}^{g},\phi,t}-\epsilon\rVert_{2}^{2}+\gamma\lVert\epsilon_{x_{t}^{g},y,t}-\epsilon_{x_{t}^{g},\phi,t}\rVert_{2}^{2}]$
$=\nabla_{\phi}\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t}^{g},\phi,t}-\epsilon\rVert_{2}^{2}-\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}^{2}+(\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}^{2}+\lVert\epsilon_{x_{t}^{r},y,t}-\epsilon\rVert_{2}^{2})+\gamma\lVert\epsilon_{x_{t}^{g},y,t}-\epsilon_{x_{t}^{g},\phi,t}\rVert_{2}^{2}]$
$=\nabla_{\phi}\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t}^{g},\phi,t}-\epsilon\rVert_{2}^{2}-\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}^{2}+\frac{1}{2}(1^{2}+1^{2})(\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}^{2}+\lVert\epsilon_{x_{t}^{r},y,t}-\epsilon\rVert_{2}^{2})+\gamma\lVert\epsilon_{x_{t}^{g},y,t}-\epsilon_{x_{t}^{g},\phi,t}\rVert_{2}^{2}]$
$\geq\nabla_{\phi}\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t}^{g},\phi,t}-\epsilon\rVert_{2}^{2}-\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}^{2}+\frac{1}{2}(\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}+\lVert\epsilon_{x_{t}^{r},y,t}-\epsilon\rVert_{2})^{2}+\gamma\lVert\epsilon_{x_{t}^{g},y,t}-\epsilon_{x_{t}^{g},\phi,t}\rVert_{2}^{2}]$
$\geq\nabla_{\phi}\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t}^{g},\phi,t}-\epsilon\rVert_{2}^{2}-\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}^{2}+\frac{1}{2}(\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon+\epsilon-\epsilon_{x_{t}^{r},y,t}\rVert_{2})^{2}+\gamma\lVert\epsilon_{x_{t}^{g},y,t}-\epsilon_{x_{t}^{g},\phi,t}\rVert_{2}^{2}]$
$=\nabla_{\phi}\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t}^{g},\phi,t}-\epsilon\rVert_{2}^{2}-\lVert\epsilon_{x_{t}^{r},\phi,t}-\epsilon\rVert_{2}^{2}+\frac{1}{2}\lVert\epsilon_{x_{t}^{r},y,t}-\epsilon_{x_{t}^{r},\phi,t}\rVert_{2}^{2}+\gamma\lVert\epsilon_{x_{t}^{g},y,t}-\epsilon_{x_{t}^{g},\phi,t}\rVert_{2}^{2}]=\nabla_{\phi}\mathcal{L}_{D}$ (17)
In practice, we find $\eta=1/2$, $\gamma\in[-1,0)$ works for most cases.
Appendix C Explanation of the workflow
To get a clearer picture of the optimization workflow, we further provide an algorithm description in Algorithm 1.
Algorithm 1 Adversarial Score Distillation for 3D tasks
Input: Prompt $y$, optimizable textual-inversion embedding or LoRA $\phi$, and pretrained diffusion model.
1:  repeat
2:     Sample a camera pose $c$.
3:     Differentiable render the 3D structure $\theta$ at pose $c$ to get a 2D image $x_{0}=g(\theta,c)$.
4:     Random select a timestep $t$ and add random noise $\epsilon$ to get $x_{t}^{g}=\alpha_{t}x_{0}+\sigma_{t}\epsilon$.
5:     Predict noise $\epsilon_{x_{t}^{g};\phi,t}$, $\epsilon_{x_{t}^{g};y,t}$ and $\epsilon_{x_{t}^{g};t}$ using the pretrained diffusion model.
6:     (Update G) Update $\theta$ with the gradient $\nabla_{\theta}\mathcal{L}_{G}^{\prime}:\mathbb{E}_{t,\epsilon}[\omega(t)(\epsilon_{x_{t}^{g};t}+\lambda(\epsilon_{x_{t}^{g};y,t}-\epsilon_{x_{t}^{g};t})-\epsilon_{x_{t}^{g};\phi,t})\frac{\partial x_{0}^{g}}{\partial\theta}]$
7:     (Update D) Update $\phi$ with the gradient $\nabla_{\phi}\mathcal{L}_{D}^{\prime}:\nabla_{\phi}\mathbb{E}_{t,\epsilon}[\lVert\epsilon_{x_{t}^{g};\phi,t}-\epsilon\rVert_{2}^{2}+\gamma\lVert\epsilon_{x_{t}^{g};y,t}-\epsilon_{x_{t}^{g};\phi,t}\rVert_{2}^{2}]$
8:  until converged
Output: 3D structure $\theta$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 12: Multi-view visualizations of generated 3D results.